OpenStudy (gabby1214):

Can anyone help solve this equation please? x÷9=3

OpenStudy (smilie2015):

x/9=3 multiplying both sides with 9 x=27

OpenStudy (osgood):

9*3=27 you use revers math for this problem 27/9=3

OpenStudy (calculusxy):

\[\large (9)\frac{ x }{ 9} = 3(9) \rightarrow x = 27\]

OpenStudy (animelover91804):

3 x 9 = 27 so... x would be 27